If three men take 3 days to complete a piece of work, how many days will 2 men take to complete the same piece of work?
Answer by reynard2007(52)

Amount of work = (Rate per Person) * (Number of Persons) * (Number of days)


Assuming that all involved people can work the same rate as the others, we solve first for the rate (what part of the work is each one able to finish in a given amount of time) per person.

(Rate Per Person) = (Amount of Work) / ( (Number of Persons)*(number of days))


Since there is only 1 job to be done, (Amount of Work) = 1. We plug in the other values to solve for the rate per person.


(Rate per Person) = (1) / ((3)*(3))
(Rate per Person) = 1 / 9


Since the question asks for the number of days two people can finish the same job, we come up with the equation based on the original:


(Number of days) = (Amount of Work) / ((Rate per Person) * (Number of Persons))
(Number of days) = 1 / ((1/9)*2)
(Number of days) = 1 / (2/9)
(Number of days) = 9/2 = 4.5 days


Two people can finish the job in 4 1/2 days.